Jorge and Anita, married taxpayers, earn $144,500 in taxable income and $54,000 in interest from an investment in City of Heflin bonds. (Use the U.S. tax rate schedule for married filing jointly). Required: If Jorge and Anita earn an additional $107,000 of taxable income, what is their marginal tax rate on this income? What is their marginal rate if, instead, they report an additional $107,000 in deductions? (For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places.)
Answer #1
Particulars Calculations Present Scenario
Taxable Income           144,500.00 A
Calculation of Federal Tax:
Upto $ 19,050 @ 10%       1,905.00
$ 19,051 to $ 77,400 @ 12%       7,001.88
Balance $ 43,100 @ 22%     14,762.00
Total Federal Tax             23,668.88 C
Average Tax Rate 16.38% D=C/A
If there is additional income           107,000.00 E
Revised Taxable Income           251,500.00 F= A+E
Calculation of Federal Tax:
Upto $ 19,050 @ 10%       1,905.00
$ 19,051 to $ 77,400 @ 12%       7,001.88
$ 77,401 to $ 165,000 @ 22%     19,271.78
Balance $ 86,500 @ 24%     20,760.00
Total Federal Tax             48,938.66 G
Differential Federal Tax             25,269.78 H=G-C
Marginal Tax Rate 23.62% I=H/E
If there is additional deduction           107,000.00 J
Revised Taxable Income             37,500.00 K=A-K
Calculation of Federal Tax:
Upto $ 19,050 @ 10%       1,905.00
Balance $ 18,450 @ 12%       2,214.00
Total Federal Tax               4,119.00 L
Differential Federal Tax           (19,549.88) M=L-C
Marginal Tax Rate 18.27% N=M/(-J)